OpenStudy (anonymous):

(X^k+1/X)^6 ;at "K" is positive integer ;find value of"K' such that the expansion has a term free of X

OpenStudy (anonymous):

@mukushla i need the value of k that make the expansion has term free of X

OpenStudy (anonymous):

if u want to find all values of \(k\) just simplify the \[(x^k+\frac{1}{x})^6\]using binomial theorem

OpenStudy (anonymous):

can u do that?

OpenStudy (anonymous):

of course but at Tr+1=6Cr+(1/x)^r +(x^K)^6-r;then Tr+1=6cr (X^6k-kr-r) at term free of "X" then ; 6k-kr-r=0 and i can not complete can you help

OpenStudy (anonymous):

very well now u need to solve \(6k-kr-r=0\) for \(r \in \text{{0,1,2,3,4,5,6}}\)

OpenStudy (anonymous):

because if \(6k-kr-r=0\) then exponent of \(x\) will be zero and that term becomes \(x\) free

OpenStudy (anonymous):

\[r \in \left\{ 0,1,2,3,4,5,6 \right\}\] becouse "r"lessthan or equal "n" and N=6 is that right?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw :)